Inessential Refrigerator Restocking

(Revision: 1.0 )

Dave Cho (dcctdw@athena.mit.edu)

October 13, 1992

Contents

1 Introduction: What is a refrigerator? . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Getting Started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3 What to Start With . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 4 Processing Your Discovery . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 5 Wait! Did you understand that? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 6 One last note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 7 Where to Get More Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 8 Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1 Introduction: What is a refrigerator?

re.frig.er.a.tor ri-’frij-*-.ra-t-*r n : something that refrigerates or keeps cool : : a cabinet or room for keeping food or other items cool : an apparatus for rapidly cooling heated liquids or vapors in a distilling process

The SIPB owns a refrigerator for the express purpose of cooling sodas and juices. It is usually appreciated by people that their sodas and juice drinks are cold.

2 Getting Started

When you see that the refrigerator is low on a particular item (for instance, Coca-Cola, also known as Coke), then you should pause and think. We realize this is difficult for some people.

3 What to Start With

First, inspect the refrigerator and determine that the refrigerator really is low. Note that low means that it has less than a reasonable number of cans left. A reasonable number is often defined, in this case, to be around 5. If the refrigerator indeed has more cans than aforementioned reasonable number, then you can skip this document, this time.

Otherwise, please continue.

4 Processing Your Discovery

At this point, you should inspect the stack of sodas and juices located next to the music equipment. If, for instance, the refrigerator is low on Coke and you see a case of Coke sitting next to the music equipment rack, then please take off any plastic covering and insert said case into the refrigerator.

5 Wait! Did you understand that?

Let’s review, just to make sure you understand. If you notice the refrigerator is low on a particular soda or juice, then (as the buzzwords in the trade go), restock the refrigerator. This means you will have to pause and do something about the empty refrigerator. (On faster platforms, there may be a noticeable increase in performance.)

6 One last note

If there are plastic rings on the sodas or juices, it is generally considered rather social to either break or slice them open with a knife, since rather nasty things can happen to various fuzzy creatures in the wild when the rings are dumped. In fact, there are some very social people who take each can out of the ring, and then slice the rings, after inserting the case into the refrigerator. This, however, should only be done by experts, and the reader who is just starting to learn from this document should not attempt such a brave move.

7 Where to Get More Information

Try calling your mother, and asking her about common courtesy. She’ll probably be happy to explain.

8 Acknowledgments

The author would like to thank Stoopidicus Maximus and Cluless Dumfuck, whose existence necessitated the creation this document.